Let F be the ring of all functions from R to R w.r.t. pointwise, addition and

multiplication. Let S be the set of all differentiable functions in F. Check whether S

is

i) a subring of F,

ii) an ideal of F.


1
Expert's answer
2021-06-17T12:40:02-0400

i) Let us check whether "S" is a subring of "F". Let "f(x),g(x)\in S," that is "f(x),g(x)" are differentiable functions. Then the function "f(x)-g(x)" is also differentiable and "(f(x)-g(x))'=f'(x)-g'(x)," and hence "f(x)-g(x)\in S." Also the function "f(x)\cdot g(x)" is differentiable and "(f(x)\cdot g(x))'=f(x)g'(x)+f'(x)g(x)," and hence "f(x)\cdot g(x)\in S." Therefore, we conclude that "S" is a subring of "F."


ii) Let us check whether "S" is an ideal of "F". Let "f(x)\notin S," for example "f(x)=|x|." Taking into account that the constant function "g:\R\to\R,\ g(x)=1," is differentiable, that is "g(x)\in S," but "g(x)\cdot f(x)=f(x)\notin S," we conclude that "S" is not an ideal of "F".